You invest 5,000 in an account at 5.5% per year simple interest. How much will you have in the account after 5 years? Round your answer to the nearest whole dollar.
Answer by stanbon(75887)

You invest 5,000 in an account at 5.5% per year simple interest. How much will you have in the account after 5 years? Round your answer to the nearest whole dollar.
--------
A(t) = P(1 + rt)
--------
A(5) = 5000(1 + 0.055*5) = 5000(1.275) = $6375
------------------
Cheers,
Stan H.
----------------------------------